# BanditSpec: Adaptive Speculative Decoding via Bandit Algorithms

### v1

@@ 5 Modeling Tokens Adversarially
The bandits problem with this assumption is often referred to as the oblivious adversarial bandits in the online learning works (Auer et al., 2002a; Lattimore & Szepesvári, 2020). It admits more general and practical setups compared to the stochastic MAB. We find the greedy decoding strategy aligns more closely to this setup in the sense that the generated tokens by the models are (potentially) fixed given the initial prompt. Hence, we present our result under the greedy decoding strategy in this section. [^2] Our analysis can also be extended to cover the sampling decoding strategy (see Remark D.3). Given a prompt $\mathrm{pt}\in\mathcal{X}^{*}$ and a bandits configuration $\nu$, the stopping time regret (1) of an arm selection algorithm $\mathtt{ALG}$ becomes
@@
[…]
@@ 6 Experiments
In this section, we conduct two sets of experiments to demonstrate the efficacy of the proposed bandit framework BanditSpec, along with UCBSpec and EXP3Spec. In the first experiment, the candidate hyperparameters are different draft models. In the second experiment, the candidate hyperparameters are different speculation lengths, where real-life LLM serving scenarios are simulated with diverse input prompts. Additional experimental results on memory utilization and additional experiments on larger models and different hardwares are provided in Appendix G. The code is accessible via https://github.com/sail-sg/BanditSpec.
@@ 6.1 Experiment with Draft Models
Experimental Setups We adopt the open-sourced LLaMA3-8B-Instruct (Dubey et al., 2024) and Qwen2-7B-Instruct (Yang et al., 2024) as the target models. The commonly-used existing speculative decoding methods PLD (Saxena, 2023), Rest (He et al., 2024), Suffix Tree (Oliaro et al., 2024; Hu et al., 2024) and Eagle-2 (Li et al., 2024b) are adopted as the baselines. Among these baselines, PLD, Rest, and Suffix Tree represent the non-parametric (or model-free) speculative decoding methods, whereas Eagle-2 represents the speculative decoding methods that utilize smaller draft models. Each of these methods corresponds to an arm in our problem.
The experiments are carried out on Spec Bench (Xia et al., 2024), Alpaca (Taori et al., 2023), Code Editor (Guo et al., 2024) and Debug Bench (Tian et al., 2024). Among these benchmarks, Spec Bench and Alpaca encompass multiple topics, while Code Editor and Debug Bench focus on coding tasks, a representative scenario for specialized models.
We record the number of accepted tokens for each speculative decoding step, as well as the wall-time for generating each complete response. The Mean Accepted Tokens (MAT) and the throughput (Tokens/s) are computed. These two metrics are widely adopted in the speculative decoding community and are positively correlated (Xia et al., 2024). In particular, Tokens/s measures the actual latency during decoding. The experiments are conducted on a single A100 and set batch size as $1$.
@@
Experimental Results We report the results of our experiments in Table 1. The proposed adaptive speculative decoding framework BanditSpec exhibits superior performance compared to existing methods in the datasets we consider. In particular, the best performance measured by Token/s is always achieved by the proposed framework. We note that although the non-parametric methods are worse than Eagle-2 in average, they are effective on a portion of prompts. Our proposed methods, UCBSpec and EXP3Spec, automatically adapt to different prompts, i.e., suffering from a small stopping time regret on each prompt. Thus, they achieve better performance than all the methods that only use a fixed model. On Debug Bench, UCBSpec can even achieve improvements of $13\%$ for LLaMA3 and $19\%$ for Qwen2. Moreover, as UCBSpec demonstrates better performance under almost all benchmarks with both two target models, this suggests that speculative decoding in real-life environments tends to be closer to the stochastic (Assumption 4.1) compared to the adversarial reward case (Assumption 5.1).
@@ 6.2 Experiment with Speculation Lengths
Experimental Setups In addition to improving the latency when batch size is $1$, our proposed algorithms also improve the throughput in real LLM serving scenarios with different batch sizes. In practical serving environments, speculative decoding does not always yield performance gains due to variations in batch size and acceptance rate. As the batch size increases, the system rapidly becomes compute-bound, while a lower acceptance rate can lead to wasted computation resources of GPU. Additionally, the execution time of the draft model contributes to an overall decrease in throughput. Given these confoundingly interrelated factors, along with latent variables such as the acceptance rate (which is unknown before verification and depends on the input prompts), we adopt a bandit-based approach to model the current throughput as the reward. Specifically, we employ UCBSpec to dynamically adjust the hyperparameter $\gamma$, the speculation length, to maximize the throughput, i.e., the number of generated tokens per second. We set the maximum speculation length $L$ as $4$, and $\gamma$ takes values in $\{0,\ldots,4\}$ where $\gamma=0$ corresponds to the canonical decoding (Algorithm 1). As the first experiment suggests UCBSpec is more in line with the real-life speculative decoding environment than EXP3Spec, we only evaluate UCBSpec in this experiment. The experiments are conducted on a single A100.
@@
Specifically, we use LLaMA3-8B-Instruct and Qwen2-7B-Instruct as the target models and adopt Eagle-1 (Li et al., 2024a), the current state-of-the-art model, as the draft model. We do not use Eagle-2 (Li et al., 2024b) because it does not support batch inference. For evaluation, we adopt Alpaca (Taori et al., 2023) as the test set, as it covers various topics, thereby simulating a realistic setting with diverse acceptance rates. To approximate real-world conditions, we randomly sample prompts from the test set to form a batch for inference, with batch sizes ranging from $1$ to $50$. As our evaluation metric, we measure the throughput improvement relative to the canonical decoding (non-speculative) baseline. Our result is averaged over $16$ independent runs to smoothen the hardware-dependent factors.
@@
Experimental Results The results are presented in Figure 3, where we reorder the $500$ sample indices in ascending order of the performance of the best hyperparameter (blue line) for easy comparison. Otherwise, the lines in this figure will not be largely monotonic. Here the “worst” and “best” lines are calculated among results of $\gamma\in\{1,\cdots,4\}$ in hindsight. Thus, we call the “best” line as the oracle best. Firstly, since the optimal hyperparameter $\gamma$ varies with different input prompts for either target model, fixing a single hyperparameter is suboptimal, e.g., in Figure 3 (b), the best hyperparameter changes from $\gamma=1$ (light green) to $\gamma=2$ (green) at sample index around $80$; and the original Eagle-1 (Li et al., 2024a) ($\gamma=4$ in purple) is even inferior to the canonical decoding ($\gamma=0$ in grey) for sample indices less than $80$. This necessitates the use of adaptive hyperparameter selection. Next, UCBSpec demonstrates competitive throughput performance, outperforming the second-best hyperparameter in most cases and closely approaching the (varying) oracle best across experiments. These benefits are obtained thanks to the adaptivity of BanditSpec.
@@ 7 Conclusions and Discussions
In this work, we propose a MAB framework together with two hyperparameter selection algorithms that adaptively choose appropriate hyperparameters to accelerate LLM inference under realistic assumptions. Both theoretical guarantees and extensive experiments are provided to demonstrate that adaptive speculative decoding via bandit algorithms can boost the performance of existing methods in a training-free manner. For future work, we would like to point some directions, improving the performance of the current algorithms.
@@
Structured Bandits Our current framework is based on the standard $K$-armed bandit model. However, broader classes of bandit problems with additional structures—such as linear bandits (Abbasi-yadkori et al., 2011) and Lipschitz bandits (Magureanu et al., 2014)—can also be considered. This aligns more closely with practical scenarios, where the number of hyperparameters can be large, and the value of $K$ may be very high when modeling the problem as a $K$-armed MAB. By leveraging such structures in MABs, we can expect to identify better hyperparameters more efficiently, thereby further accelerating the optimization process.
Robust bandits and Non-stationary bandits As indicated by the experimental result, the real-life speculative decoding environment is closer to the stochastic reward case (Assumption 4.1) than the adversarial reward case (Assumption 5.1). Therefore, one direction for future work is to consider the settings “in between”, e.g., robust bandits in the presence of adversarial corruptions (Ding et al., 2022; Zhong et al., 2021), or non-stationary bandits (Cao et al., 2019; Besbes et al., 2014; Hou et al., 2024) where the mean number of accepted tokens can vary across time. These settings are more benign than the adversarial reward assumption and can be exploited to accelerate the inference.
Contextual Bandits Another direction is to explore contextual bandits, where the environment reveals additional information that can be leveraged to reduce the learning burden (Luo et al., 2018; Kato & Ariu, 2021).
@@ Appendix A Related Works
Speculative Decoding Speculative decoding is proposed in Leviathan et al. (2023); Chen et al. (2023), where the draft model only generates a single chain of draft tokens. Then a line of works extends the chain structure to the tree structure (Miao et al., 2023; Cai et al., 2024; Du et al., 2024; Li et al., 2024a; Hu & Huang, 2024). In these works, the draft tokens are organized as a connected tree. To further improve the number of accepted tokens, previous works propose to generate tokens in a batch manner, i.e., the draft tokens are organized as multiple disconnected parts. SpecTr (Sun et al., 2024c) views this problem from the optimal transport perspective and derives the algorithm that is optimal up to a multiplicative constant. Khisti et al. (2025) derives the canonical form of this problem and design the relaxed optimization algorithms. All these algorithms verify the draft tokens in a token by token manner. Sun et al. (2024b) proposes to verify all the draft tokens as a whole block, which further boosts the acceleration ratio. Sun et al. (2024a) proposes to fit the speculative deciding into a hierarchical structure, which multiple draft models with various sizes are generating and verifying tokens. The smaller model will generate more tokens. This fine-grained behavior improve the overall performance of the system. Liu et al. (2023) design algorithms to update the draft model parameters in an online manner, which makes the draft model adaptive to the current context. Liu et al. (2024) and Huang et al. (2024) aim to optimize the speculative length in a training and training-free manner (more discussions on SpecDec++ (Huang et al., 2024) are provided in Appendix B.4). Chen et al. (2024) optimizes the hyperparameters related to the hardware by dynamic programming in an offline manner. We also note that there are a series of non-parametric speculative decoding algorithms (Hu et al., 2024; Oliaro et al., 2024), i.e., the draft model itself does not require any training procedures. Yin et al. (2024) derives the theoretical analysis of the speculative decoding.
@@
Multi-Armed Bandit The multi-armed bandit problem is a fundamental topic in decision theory and reinforcement learning, with various algorithms developed to address the exploration-exploitation trade-off. The standard stochastic $K$-armed bandit problem is firstly introduced by Robbins (1952) and then studied by Lai & Robbins (1985). There has been a major theoretical advancement with the introduction of Upper Confidence Bound (UCB) algorithms (Auer et al., 2002b). Various algorithms have been proposed to achieve improved theoretical guarantees and practical performance since then (Garivier & Cappé, 2011; Bubeck et al., 2012). Beyond UCB-type algorithms, sampling-based algorithms, such as Thompson Sampling (Agrawal & Goyal, 2012, 2017; Russo et al., 2017) and sampling via bootstrap (Kveton et al., 2019; Wan et al., 2023), have also exhibited strong empirical performance with provable regret bounds. Furthermore, the problem has been extended to the adversarial settings where the rewards are no longer stochastic (Auer et al., 2002a; Bubeck et al., 2012). We refer to Lattimore & Szepesvári (2020) for a comprehensive introduction of the Multi-Armed Bandit problem.
@@ B.2 Discussion on UCBSpec
In contrast, Thompson Sampling (Agrawal & Goyal, 2012) and KL-UCB (Garivier & Cappé, 2011) generally achieve better empirical regret bounds than UCB1 (Auer et al., 2002b). However, Thompson Sampling requires maintaining the posterior distribution and sampling from it to select the arm to pull; whereas KL-UCB involves solving an optimization problem for arm selection. These steps add additional complexity to the algorithms.
@@ B.3 Discussion on Adaptive Adversary
We consider the oblivious adversary in this paper where the numbers of accepted tokens generated by all hyperparameters at all time steps, i.e., $\{y_{i,t}\}_{i\in[K],t\in\mathbb{N}}$, are fixed before the decoding process starts. One may be interested in considering the adapative adversary, where the environment (adversary) can choose the number of accepted tokens generated by $S_{I_{t}}$ based on (part of) the history information $\mathcal{H}_{t-1}$ and $I_{t}$ (Arora et al., 2012). This adversary is more malicious than the oblivious one and the regret is expected to be even larger than the current one in Theorem 5.3. As our empirical experiments suggest that the practical scenario aligns more closely with the stochastic reward assumption (Assumption 4.1) and deviates from the oblivious adversarial reward assumption (Assumption 5.1), we believe it is unnecessary to consider the adaptive adversarial reward.
B.4 Discussion on SpecDec++ (Huang et al., 2024)
We compare the proposed methods with an existing adaptive speculative decoding method, SpecDec++ (Huang et al., 2024), in this section.
SpecDec++ (Huang et al., 2024) is adaptive in choosing the speculation length, achieving good performance compared to the vanilla speculative decoding method (Leviathan et al., 2023; Chen et al., 2023). It trains an acceptance probability prediction head and stops drafting new tokens when the predicted rejection probability reaches certain threshold.

### v2

@@ 5 Modeling Tokens Adversarially
The bandit problem with this assumption is often referred to as the oblivious adversarial bandits in the online learning works (auer2002nonstochastic; lattimore2020bandit). It admits more general and practical setups compared to the stochastic MAB. We find the greedy decoding strategy aligns more closely to this setup in the sense that the generated tokens by the models are (potentially) fixed given the initial prompt. Hence, we present our result under the greedy decoding strategy in this section. [^2] Our analysis can also be extended to cover the sampling decoding strategy (see Remark D.3). Given a prompt $\mathrm{pt}\in\mathcal{X}^{*}$ and a bandits configuration $\nu$, the stopping time regret (1) of an arm selection algorithm $\mathtt{ALG}$ becomes
@@
Algorithm Design We present our arm selection algorithm in Algorithm 5, which is an abridged version of the full version $\textsc{BanditSpec}(\textsc{EXP3Spec})$ delineated in Algorithm 9. This algorithm modifies the anytime EXP3 algorithm (lattimore2020bandit) to suit the speculative decoding application. In terms of the algorithm design, the main difference lies in the change of the stopping rule. We highlight that while the stopping time of the algorithm is random, the anytime feature of Algorithm 5 does not require any information about the time horizon. This is achieved by the vanishing sequence of learning rates $\{\eta_{t}\}_{t\in\mathbb{N}}$ which can be elegantly adapted to the unknown stopping time. With regard to the analysis, previous works (auer2002nonstochastic; bubeck2012regret; lattimore2020bandit) only consider the gap between the cumulated rewards over the same fixed horizon $T$, i.e., $\max_{i\in[K]}\mathbb{E}[\sum_{t\in[T]}y_{i,t}-y_{I_{t},t}]$. In contrast, we need to upper bound the stopping time regret in (12) where the baseline $\mathtt{ALG}_{i}$ and any proposed algorithm $\mathtt{ALG}$ have different termination times in general. Thus, the analysis is much more involved.
@@ 6 Experiments
In this section, we conduct two sets of experiments to demonstrate the efficacy of the proposed bandit framework BanditSpec, along with UCBSpec and EXP3Spec. In the first experiment, the candidate hyperparameters are different draft models. In the second experiment, the candidate hyperparameters are different speculation lengths, where real-life LLM serving scenarios are simulated with diverse input prompts. Additional experimental results on memory utilization and additional experiments on larger models and different hardware are provided in Appendix G. The code is accessible via https://github.com/sail-sg/BanditSpec.
@@ 6.1 Experiment with Draft Models
Experimental Setups We adopt the open-sourced LLaMA3-8B-Instruct (grattafiori2024llama3herdmodels) and Qwen2-7B-Instruct (yang2024qwen2technicalreport) as the target models. The commonly-used existing speculative decoding methods PLD (saxena2023prompt), Rest (he2024rest), Suffix Tree (oliaro2024suffixdecoding; hu2024sam) and Eagle-2 (li2024eagle2) are adopted as the baselines. Among these baselines, PLD, Rest, and Suffix Tree represent the non-parametric (or model-free) speculative decoding methods, whereas Eagle-2 represents the speculative decoding methods that utilize smaller draft models. Each of these methods corresponds to an arm in our problem.
The experiments are carried out on Spec Bench (xia2024unlocking), Alpaca (Taori2023alpaca), Code Editor (guo2024codeeditorbench) and Debug Bench (tian2024debugbench). Among these benchmarks, Spec Bench and Alpaca encompass multiple topics, while Code Editor and Debug Bench focus on coding tasks, a representative scenario for specialized models.
We record the number of accepted tokens for each speculative decoding step, as well as the wall-time for generating each complete response. The Mean Accepted Tokens (MAT) and the throughput (Tokens/s) is computed. These two metrics are widely adopted in the speculative decoding community and are positively correlated (xia2024unlocking). In particular, Tokens/s measures the actual latency during decoding. The experiments are conducted on a single A100 and set the batch size to $1$.
@@
Experimental Results We report the results of our experiments in Table 1. The proposed adaptive speculative decoding framework BanditSpec exhibits superior performance compared to existing methods in the datasets we consider. In particular, the best performance measured by Token/s is always achieved by the proposed framework. We note that although the non-parametric methods are worse than Eagle-2 in average, they are effective on a portion of prompts. Our proposed methods, UCBSpec and EXP3Spec, automatically adapt to different prompts, i.e., suffering from a small stopping time regret on each prompt. Thus, they achieve better performance than all the methods that only use a fixed model. On Debug Bench, UCBSpec can even achieve improvements of $13\%$ for LLaMA3 and $19\%$ for Qwen2. Moreover, as UCBSpec demonstrates better performance under almost all benchmarks with both target models, this suggests that speculative decoding in real-life environments tends to be closer to the stochastic (Assumption 4.1) compared to the adversarial reward case (Assumption 5.1).
@@ 6.2 Experiment with Speculation Lengths
Experimental Setups In addition to improving the latency when batch size is $1$, our proposed algorithms also improve the throughput in real LLM serving scenarios with different batch sizes. In practical serving environments, speculative decoding does not always yield performance gains due to variations in batch size and acceptance rate. As the batch size increases, the system rapidly becomes compute-bound, while a lower acceptance rate can lead to wasted computation resources of the GPU. Additionally, the execution time of the draft model contributes to an overall decrease in throughput. Given these confoundingly interrelated factors, along with latent variables such as the acceptance rate (which is unknown before verification and depends on the input prompts), we adopt a bandit-based approach to model the current throughput as the reward. Specifically, we employ UCBSpec to dynamically adjust the hyperparameter $\gamma$, the speculation length, to maximize the throughput, i.e., the number of generated tokens per second. We set the maximum speculation length $L$ as $4$, and $\gamma$ takes values in $\{0,\ldots,4\}$ where $\gamma=0$ corresponds to the canonical decoding (Algorithm 1). As the first experiment suggests UCBSpec is more in line with the real-life speculative decoding environment than EXP3Spec, we only evaluate UCBSpec in this experiment. The experiments are conducted on a single A100.
@@
Specifically, we use LLaMA3-8B-Instruct and Qwen2-7B-Instruct as the target models and adopt Eagle-1 (li2024eagle), the current state-of-the-art model, as the draft model. We do not use Eagle-2 (li2024eagle2) because it does not support batch inference. For evaluation, we adopt Alpaca (Taori2023alpaca) as the test set, as it covers various topics, thereby simulating a realistic setting with diverse acceptance rates. To approximate real-world conditions, we randomly sample prompts from the test set to form a batch for inference, with batch sizes ranging from $1$ to $50$. As our evaluation metric, we measure the throughput improvement relative to the canonical decoding (non-speculative) baseline. Our result is averaged over $16$ independent runs to smooth the hardware-dependent factors.
@@
Experimental Results The results are presented in Figure 3, where we reorder the $500$ sample indices in ascending order of the performance of the best hyperparameter (blue line) for easy comparison. Otherwise, the lines in this figure will not be largely monotonic. Here the “worst” and “best” lines are calculated among results of $\gamma\in\{1,\cdots,4\}$ in hindsight. Thus, we call the “best” line as the oracle best. Firstly, since the optimal hyperparameter $\gamma$ varies with different input prompts for either target model, fixing a single hyperparameter is suboptimal, e.g., in Figure 3 (b), the best hyperparameter changes from $\gamma=1$ (light green) to $\gamma=2$ (green) at a sample index around $80$; and the original Eagle-1 (li2024eagle) ($\gamma=4$ in purple) is even inferior to the canonical decoding ($\gamma=0$ in grey) for sample indices less than $80$. This necessitates the use of adaptive hyperparameter selection. Next, UCBSpec demonstrates competitive throughput performance, outperforming the second-best hyperparameter in most cases and closely approaching the (varying) oracle best across experiments. These benefits are obtained thanks to the adaptivity of BanditSpec.
@@ 7 Conclusions and Discussions
In this work, we propose a MAB framework together with two hyperparameter selection algorithms that adaptively choose appropriate hyperparameters to accelerate LLM inference under realistic assumptions. Both theoretical guarantees and extensive experiments are provided to demonstrate that adaptive speculative decoding via bandit algorithms can boost the performance of existing methods in a training-free manner. For future work, we would like to point out some directions, improving the performance of the current algorithms.
@@
Structured Bandits Our current framework is based on the standard $K$-armed bandit model. However, broader classes of bandit problems with additional structures—such as linear bandits (Abbasi2011improved) and Lipschitz bandits (magureanu2014lipschitz)—can also be considered. This aligns more closely with practical scenarios, where the number of hyperparameters can be large, and the value of $K$ may be very high when modeling the problem as a $K$-armed MAB. By leveraging such structures in MABs, we can expect to identify better hyperparameters more efficiently, thereby further accelerating the optimization process.
Robust bandits and Non-stationary bandits As indicated by the experimental result, the real-life speculative decoding environment is closer to the stochastic reward case (Assumption 4.1) than the adversarial reward case (Assumption 5.1). Therefore, one direction for future work is to consider the settings “in between”, e.g., robust bandits in the presence of adversarial corruptions (Ding2022Robust; zhong2021probabilistic), or non-stationary bandits (cao2019nearly; Besbes2014Stochastic; hou2024almost) where the mean number of accepted tokens can vary across time. These settings are more benign than the adversarial reward assumption and can be exploited to accelerate the inference.
Contextual Bandits Another direction is to explore contextual bandits, where the environment reveals additional information that can be leveraged to reduce the learning burden (luo2018contextual; kato2021role).
@@ Appendix A Related Works
Speculative Decoding Speculative decoding is proposed in leviathan2023fast; chen2023accelerating, where the draft model only generates a single chain of draft tokens. Then a line of works extends the chain structure to the tree structure (miao2023SpecInferAG; cai2024medusa; du2024glide; li2024eagle; huaccelerated). In these works, the draft tokens are organized as a connected tree. To further improve the number of accepted tokens, previous works propose to generate tokens in a batch manner, i.e., the draft tokens are organized as multiple disconnected parts. SpecTr (sun2024spectr) views this problem from the optimal transport perspective and derives the algorithm that is optimal up to a multiplicative constant. khisti2024multi derives the canonical form of this problem and designs the relaxed optimization algorithms. All these algorithms verify the draft tokens in a token-by-token manner. sun2024block proposes to verify all the draft tokens as a whole block, which further boosts the acceleration ratio. sun2024triforce proposes to fit the speculative deciding into a hierarchical structure, where multiple draft models with various sizes are generating and verifying tokens. The smaller model will generate more tokens. This fine-grained behavior improves the overall performance of the system. liu2023OnlineSD design algorithms to update the draft model parameters in an online manner, which makes the draft model adaptive to the current context. liu2024optimizing and huang2024specdec++ aim to optimize the speculative length in a training and training-free manner (more discussions on SpecDec++ (huang2024specdec++) are provided in Appendix B.4). chen2024SequoiaSR optimizes the hyperparameters related to the hardware by dynamic programming in an offline manner. We also note that there is a series of non-parametric speculative decoding algorithms (hu2024sam; oliaro2024suffixdecoding), i.e., the draft model itself does not require any training procedures. yin2024theoretical derives the theoretical analysis of the speculative decoding.
@@
Multi-Armed Bandit The multi-armed bandit problem is a fundamental topic in decision theory and reinforcement learning, with various algorithms developed to address the exploration-exploitation trade-off. The standard stochastic $K$-armed bandit problem was first introduced by robbins1952some and then studied by lai1985asymptotically. There has been a major theoretical advancement with the introduction of Upper Confidence Bound (UCB) algorithms (auer2002finite). Various algorithms have been proposed to achieve improved theoretical guarantees and practical performance since then (garivier2011KLUCB; bubeck2012regret). Beyond UCB-type algorithms, sampling-based algorithms, such as Thompson Sampling (agrawal2012analysis; agrawal2017near; russo2017ATO) and sampling via bootstrap (kveton2019garbage; wan2023multiplier), have also exhibited strong empirical performance with provable regret bounds. Furthermore, the problem has been extended to the adversarial settings where the rewards are no longer stochastic (auer2002nonstochastic; bubeck2012regret). We refer to lattimore2020bandit for a comprehensive introduction to the Multi-Armed Bandit problem.
@@ B.2 Discussion on UCBSpec
In contrast, Thompson Sampling (agrawal2012analysis) and KL-UCB (garivier2011KLUCB) generally achieve better empirical regret bounds than UCB1 (auer2002finite). However, Thompson Sampling requires maintaining the posterior distribution and sampling from it to select the arm to pull; whereas KL-UCB involves solving an optimization problem for arm selection. These steps add additional complexity to the algorithms.
@@ B.3 Discussion on Adaptive Adversary
We consider the oblivious adversary in this paper where the number of accepted tokens generated by all hyperparameters at all time steps, i.e., $\{y_{i,t}\}_{i\in[K],t\in\mathbb{N}}$, is fixed before the decoding process starts. One may be interested in considering the adapative adversary, where the environment (adversary) can choose the number of accepted tokens generated by $S_{I_{t}}$ based on (part of) the history information $\mathcal{H}_{t-1}$ and $I_{t}$ (arora2012online). This adversary is more malicious than the oblivious one and the regret is expected to be even larger than the current one in Theorem 5.3. As our empirical experiments suggest that the practical scenario aligns more closely with the stochastic reward assumption (Assumption 4.1) and deviates from the oblivious adversarial reward assumption (Assumption 5.1), we believe it is unnecessary to consider the adaptive adversarial reward.
B.4 Discussion on SpecDec++ (huang2024specdec++)
We compare the proposed methods with an existing adaptive speculative decoding method, SpecDec++ (huang2024specdec++), in this section.
SpecDec++ (huang2024specdec++) is adaptive in choosing the speculation length, achieving good performance compared to the vanilla speculative decoding method (leviathan2023fast; chen2023accelerating). It trains an acceptance probability prediction head and stops drafting new tokens when the predicted rejection probability reaches a certain threshold.